# Ghost Noise for Regularizing Deep Neural Networks

### v1

@@ 2.2 Reducing train-test discrepancy
Unbounded output range of XBN: With XBN the output for a single element ${\bm{x}}_{i}$ will be:
@@
The numerator here has no dependency on ${\bm{x}}_{i}$ and if we ignore $\varepsilon$, it can therefore be arbitrarily small in comparison, making the output unbounded. For example when all ${\bm{x}}_{j},j\neq i$ are identical, the numerator of Equation 6 will be $\sqrt{\varepsilon}$, which is close to 0 in practice (the default value in PyTorch is $\varepsilon=10^{-5}$). Large values like these can destabilize training. We observe this in practice when using smaller batch sizes, likely due to ${\bm{\sigma}}_{i}^{2}$ randomly being close to zero for some ghost batches. It may be possible to mitigate this issue, for example by clamping the computed sigma or the resulting output, but we do not pursue this further here. The preceding analysis indicates that self-dependency can be beneficial for stability but larger batch sizes are preferable to minimize the resulting train-test discrepancy. We therefore seek alternative ways of obtaining the Ghost Noise of small batches that do not rely on decreasing the normalization batch size. This would confer stability while also avoiding a significant train-test discrepancy by keeping the self-dependency small.
2.3 Modelling Ghost Batch Normalization as Double Normalization
One of our key insights is that performing a standard batch normalization before ghost batch normalization does not change the output of the GBN. For a batch $\bm{X}$, batch normalization BN and ghost batch normalization GBN we can write this as:
@@ 2.3 Modelling Ghost Batch Normalization as Double Normalization
This follows from the fact that normalization is invariant to affine transformations of the inputs, i.e. $f(\bm{X})={\bm{a}}\bm{X}+{\bm{b}}$} where ${\bm{a}},{\bm{b}}$ are constant vectors. When $\varepsilon$ is small enough to be ignored, we have:
$\text{BN}({\bm{a}}\bm{X}+{\bm{b}})=\frac{({\bm{a}}\bm{X}+{\bm{b}})-\bm{\mu}({\bm{a}}\bm{X}+{\bm{b}})}{\sqrt{{\bm{\sigma}}_{i}^{2}({\bm{a}}\bm{X}+{\bm{b}})+\varepsilon}}=\frac{{\bm{a}}\bm{X}+{\bm{b}}-{\bm{a}}\bm{\mu}(\bm{X})-{\bm{b}}}{\sqrt{{\bm{a}}^{2}{\bm{\sigma}}_{i}^{2}(\bm{X})+\varepsilon}}=\frac{\bm{X}-\bm{\mu}(\bm{X})}{\sqrt{{\bm{\sigma}}_{i}^{2}(\bm{X})+\varepsilon/{\bm{a}}^{2}}}\approx\text{BN}(\bm{X})$ (8)
where we treat $\bm{\mu}$ and $\bm{\sigma}$} as functions. For a given batch, BN is an affine transformation with:
${\bm{a}}=({\bm{\sigma}}_{i}^{2}(\bm{X})+\varepsilon)^{-\frac{1}{2}},\qquad{\bm{b}}=-({\bm{\sigma}}_{i}^{2}(\bm{X})+\varepsilon)^{-\frac{1}{2}}\bm{\mu}(\bm{X})$ (9)
The coefficients of course depend on $\bm{X}$, but that does not matter for the following normalization.
This decomposition of GBN into two successive normalization operations lets us isolate the differences between standard batch normalization and ghost batch normalization. Ignoring $\varepsilon$, the double normalization $\text{GBN}(\text{BN}(\bm{X}))$ can be formulated as:
@@
where we split the normalized batch into $g$ ghost batches and show the following GBN for the $j$-th subgroup. We can now write the $\bm{\mu}_{j}$ and ${\bm{\sigma}}_{j}$ of the original $\text{GBN}(\bm{X})$ setup in Equation 2 as:
$\bm{\mu}_{j}=\bm{\mu}+{\bm{\sigma}}\hat{\bm{\mu}}_{j},\qquad{\bm{\sigma}}_{j}={\bm{\sigma}}\hat{{\bm{\sigma}}}_{j}$ (11)
GBNwould be equivalent to BN if $\hat{\bm{\mu}}_{j}=0$ and $\hat{{\bm{\sigma}}}_{j}=1$. This happens when the ghost batches are identical, but generally they contain different samples resulting in random variations in $\hat{\bm{\mu}}_{j}$ and $\hat{{\bm{\sigma}}}_{j}$. These fluctuations cause noise with two components, additive (or shifting) noise from $\hat{\bm{\mu}}_{j}$ and multiplicative (or scaling) noise from $\hat{{\bm{\sigma}}}_{j}$. The increased train-test discrepancy of GBN comes from the dependency of $\hat{\bm{\mu}}_{j}$ and $\hat{{\bm{\sigma}}}_{j}$ on a given sample.
2.4 Ghost Noise Injection
The preceding analysis isolates $\hat{\bm{\mu}}_{j}$ and $\hat{{\bm{\sigma}}}_{j}$ in Equation 11 as the source of Ghost Noise and the train-test discrepancy from self-dependency. We propose replacing the Ghost Batch Normalization in the double normalization setup with a new method, Ghost Noise Injection (GNI) given by:
${\bm{Y}}=\frac{\hat{\bm{X}}-{\mathbf{m}}}{{\mathbf{s}}}$ (12)
where $\mathbf{m}$} and $\mathbf{s}$} are independent and identically distributed random vectors corresponding to each sample in the input $\hat{\bm{X}}$. The elements of $\mathbf{m}$} and $\mathbf{s}$} are computed as the mean and standard deviation of a randomly sampled (with replacement) subsets of size $N$ from $\hat{\bm{X}}$. We will refer to the subset size as ghost batch size like for GBN. We treat the resulting $\mathbf{m}$} and $\mathbf{s}$} as pure noise, i.e. we don’t backpropagate through their computation. GNI has two main advantages over GBN:
  • We can freely select normalization batch size $N$, and it does not need to divide the accelerator batch size used.
Wider Applicability of GNI: Ghost Noise Injection does not necessarily need to follow batch normalization. In this case we can instead perform GNI as ${\bm{Y}}=\frac{\hat{\bm{X}}-({\mathbf{m}}-\bm{\mu})}{{\mathbf{s}}/{\bm{\sigma}}}$ where $\bm{\mu}$ and $\bm{\sigma}$} are computed like in batch normalization (but without backpropagating through them). This lets us decouple the noise from the normalization, allowing us to inject ghost noise anywhere in the network.
2.5 The Distribution of Ghost Noise
The estimated mean $\hat{\bm{\mu}}_{i}$ and standard deviation $\hat{{\bm{\sigma}}}_{i}$ from the $i$-th ghost batch, can be interpreted as bootstrapped statistics derived from the empirical distribution $P_{\hat{\bm{X}}}$. Following the normalization in Equation 10, the variable $\hat{\bm{X}}$ exhibits a zero mean and unit variance. Assuming that the individual elements of $\hat{\bm{X}}$ are normally distributed allows us to derive an analytical distribution for the mean and variance, giving us additional insights into the workings of GBN. For this section we focus on the distribution for a single channel $c$ in the $g$-th ghost batch, which we denote $\hat{\mu}_{gc}$ and $\hat{\sigma}_{gc}$.
2.5.1 Fully Connected Layers
If we assume that the output of batch normalization is independent and normally distributed, the normalization statistics $\hat{\bm{\mu}}_{i}$ and $\hat{{\bm{\sigma}}}_{i}^{2}$ are computed over a sample $\hat{\bm{X}}=[\hat{\bm{x}}_{1},..,\hat{\bm{x}}_{N}]$ of $N$ variables from $\mathcal{N}(0,1)$. We can the derive the distribution of the sample mean, and therefore the shift noise, as:
$\hat{\mu}_{gc}=\frac{1}{N}\sum_{i=1}^{N}\hat{x}_{ic}\sim\mathcal{N}\left(0,\frac{1}{N}\right)$ (13)
Since the sum of standard normally distributed variables follows a chi-squared distribution we get:
$\hat{\sigma}_{gc}^{2}=\frac{1}{N}\sum_{i=1}^{N}(\hat{x}_{ic}-0)^{2}\sim\frac{1}{N}\mathcal{\chi}^{2}(N)$ (14)
Analytical Ghost Noise Injection: From the above analysis, instead of computing ghost statistics from a sampled batch, we could directly sample $\hat{\bm{\mu}}_{g}$ and $\hat{{\bm{\sigma}}}_{g}^{2}$ from the analytical distribution for each channel. Let $\hat{\bm{X}}$ be $\bm{X}$ after batch normalization like before. For each channel $c$, we can then inject noise using:
Comparison to Gaussian Dropout: Here we would like to point out the similarity between the scaling noise and Gaussian Dropout which is written as:
where $\mathbf{t}$} is sampled element-wise from the Gaussian distribution and $p$ is interpreted like the drop probability in standard dropout ($p=0$ for no dropout). Both the scaling noise and Gaussian dropout are multiplicative with similar but slightly different distributions.
2.5.2 Convolutional Layers
In convolutional neural networks the batch statistics are computed across both the batch and spatial dimensions (e.g. the height and width of an image). The data can therefor vary in two separate ways, between samples across the batch dimension, i.e. inter-sample variance and within a single sample across the spatial dimension, i.e. intra-sample variance. This differs from the fully connected case where all variance arises across the batch dimension and there is no concept of intra-sample variance.
To investigate the effects of this, we analyze a simple model. We will assume that $\bm{X}\in\mathbb{R}^{B\times C\times I}$, where $I$ is the spatial dimension e.g. $I=H\times W$ where $H$ is the height and $W$ the width for image data. We further assume the true mean of $b$-th batch in a specific channel $c$ is sampled i.i.d. from a normal distribution, i.e. $\mu_{bc}\overset{i.i.d.}{\sim}\mathcal{N}(0,\sigma_{Bc}^{2})$, where $\sigma_{Bc}^{2}$ denotes the inter-sample variance and only varies across $c$, i.e. it is constant across $B$. We model the $i$-th spatial location in the $b$-th batch as being sampled from $x_{bic}\overset{i.i.d.}{\sim}\mathcal{N}(\mu_{bc},\sigma_{Ic}^{2})$}, with an intra-sample variance $\sigma_{Ic}^{2}$ that does not vary across $I$.After batch normalization, we always have unit variance, which means $\sigma_{Bc}^{2}+\sigma_{Ic}^{2}=1$, for any $c$. Now assume we sample a random ghost batch of size $N$. As the sample mean is an average of all samples, we still have it following a normal distribution, and the mean and variance can be calculated as follows:
$\operatorname{\operatorname{\mathbb{E}}}[\hat{\mu}_{gc}]=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nic}\right]=\frac{1}{NI}\sum_{n}\sum_{i}\operatorname{\operatorname{\mathbb{E}}}[x_{nic}]=0$ (17)
$\mathrm{Var}[\hat{\mu}_{gc}]$ $=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nic}\right]^{2}$ (18)
$=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nic}-\mu_{nc}+\mu_{nc}\right]^{2}$
$=\frac{1}{N^{2}I^{2}}\sum_{n}\sum_{i}\sigma_{Ic}^{2}+\frac{1}{N^{2}I^{2}}NI^{2}\sigma_{Bc}^{2}$
$=\frac{1}{NI}\sigma_{Ic}^{2}+\frac{1}{N}\sigma_{Bc}^{2}$
The distribution of shift noise in channel $c$ can thus be written as:
For the scale noise in a specific channel $c$, we have
$\hat{\sigma}_{gc}^{2}=$ $\frac{1}{NI}\sum_{i}\sum_{n}(x_{nic}-\mu_{nc}+\mu_{nc})^{2}$ (20)
$=$ $\frac{1}{NI}\sum_{i}\sum_{n}\left((x_{nic}-\mu_{nc})^{2}+(\mu_{nc}-\mu)^{2}+2(x_{nic}-\mu_{nc})(\mu_{nc}-\mu)\right)$
@@ 2.5.2 Convolutional Layers
$\overset{\overline{x}_{nc}\approx\operatorname{\operatorname{\mathbb{E}}}x_{nc}}{\approx}$ $\frac{1}{NI}\sum_{i}\sum_{n}\sigma_{Ic}^{2}\left(\frac{x_{nic}-\mu_{nc}}{\sigma_{Ic}}\right)^{2}+\frac{1}{N}\sigma_{Bc}^{2}\sum_{n}\left(\frac{\mu_{nc}-\mu}{\sigma_{Bc}}\right)^{2}+\frac{1}{N}(\mu_{nc}-\mu_{nc})(\mu_{nc}-\mu)$
$=$ $\frac{1}{NI}\sum_{i}\sum_{n}\sigma_{Ic}^{2}\left(\frac{x_{nic}-\mu_{nc}}{\sigma_{Ic}}\right)^{2}+\frac{1}{N}\sigma_{Bc}^{2}\sum_{n}\left(\frac{\mu_{nc}-\mu}{\sigma_{Bc}}\right)^{2}$
It follows that $\hat{\sigma}_{gc}^{2}$ approximately follows a mixture of two chi-square distributions.
When $\sigma_{Ic}^{2}=0$, the results coincide with the 1d case. However, the analysis in 2D case gives us more freedom to separate inter- ($\sigma_{Bc}^{2}$) and intra-($\sigma_{Ic}^{2}$) sample variances. Another insightful aspect is that it provides more understanding of the differences across channels. As both $\sigma_{Bc}^{2}$ and $\sigma_{Ic}^{2}$ are channel-specific, there is indeed a channel-wise noise effect, which we will provide an empirical proof in Figure 3.
3 Experiments
Figure 2: CIFAR-100 ResNet-18 validation accuracy versus ghost batch size and dropout probability for different methods. Each line is the average of five runs and the shaded area shows the standard deviation. The dotted lines show the standard batch normalization performance and the maximum for ghost batch normalization. For dropout (DO, lower right panel), B=Bernoulli, G=Gaussian, E=elementwise and C=channelwise. See Section 3 for further experimental details and discussion.

### v2

@@ Appendix B Unbounded output range of XBN
Unbounded output range of XBN
With XBN the output for a single element ${\bm{x}}_{i}$ will be:
@@
The numerator here has no dependency on ${\bm{x}}_{i}$ and if we ignore $\varepsilon$, it can therefore be arbitrarily small in comparison, making the output unbounded. For example when all ${\bm{x}}_{j},j\neq i$ are identical, the numerator of equation 19 will be $\sqrt{\varepsilon}$, which is close to 0 in practice (the default value in PyTorch is $\varepsilon=10^{-5}$). Large values like these can destabilize training. We observe this in practice when using smaller batch sizes, due to ${\bm{\sigma}}_{i}^{2}$ randomly being close to zero for some ghost batches. It may be possible to mitigate this issue, for example by clamping the computed sigma or the resulting output, but we do not pursue this further here.
Appendix C Derivation of equation 10
Consider the double normalization setup described in Section 2.2, repeated below:
@@ Appendix C Derivation of equation 10
We identified the second transformation (Equation 21) as the source of the increased noise and train-test discrepancy in Ghost Batch Normalization. Our goal in Ghost Noise Injection is twofold, decoupling the noise from the normalization and eliminating the increased train-test discrepancy.
Eliminating the Normalization
We first show how we can decouple Equation 21 from Equation 20. The output $\widetilde{\bm{X}}_{g}$ in Equation 21 can be written as:
$\widetilde{{\bm{X}}}_{g}$ $=\frac{1}{\hat{{\bm{\sigma}}}_{g}}\left(\frac{{\bm{X}}_{g}-{\bm{\mu}}}{{\bm{\sigma}}}-\hat{{\bm{\mu}}}_{g}\right)$ (22)
$=\frac{{\bm{X}}_{g}}{\hat{{\bm{\sigma}}}_{g}{\bm{\sigma}}}-\frac{{\bm{\mu}}}{\hat{{\bm{\sigma}}}_{g}{\bm{\sigma}}}-\frac{\hat{{\bm{\mu}}}_{g}}{\hat{{\bm{\sigma}}}_{g}}$ (23)
We can undo the effects of the original normalization by multiplying this output by $\bm{\sigma}$} and then reverse the relative shift by adding ${\bm{\mu}}/\hat{{\bm{\sigma}}}_{g}$, accounting by the fact that in $\widetilde{{\bm{X}}}_{g}$ the original shift has been scaled by $\hat{{\bm{\sigma}}}_{g}^{-1}$. We will refer to this “denormalized” output of Equation 21 as $\text{GBN}_{\Delta}$ which we can write as:
$\text{GBN}_{\Delta}({\bm{X}}):$ $={\bm{\sigma}}\widetilde{{\bm{X}}}_{g}+\frac{{\bm{\mu}}}{\hat{{\bm{\sigma}}}_{g}}$ (24)
$=\frac{{\bm{X}}-{\bm{\sigma}}\hat{{\bm{\mu}}}_{g}}{\hat{{\bm{\sigma}}}_{g}}$ (25)
The GBN statistics $\hat{{\bm{\mu}}}_{g}$ and $\hat{{\bm{\sigma}}}_{g}$ can be expressed in terms of the original inputs $\bm{X}$}:
$\hat{{\bm{\mu}}}_{g}$ $=\frac{1}{NHW}\sum_{n,h,w}\hat{{\bm{X}}}_{g}$ (26)
$=\frac{1}{NHW}\sum_{n,h,w}\left(\frac{{\bm{X}}_{g}-{\bm{\mu}}}{{\bm{\sigma}}}\right)$ (27)
$=\frac{1}{{\bm{\sigma}}}\Big(\frac{1}{NHW}\sum_{n,h,w}({\bm{X}}_{g})-{\bm{\mu}}\Big)$ (28)
$=:\frac{{\bm{\mu}}_{g}-{\bm{\mu}}}{{\bm{\sigma}}}$ (29)
and we can similarly show that:
$\hat{{\bm{\sigma}}}_{g}^{2}$ $=\frac{1}{{\bm{\sigma}}^{2}}\frac{1}{NHW}\sum_{n,h,w}({\bm{X}}_{g}-{\bm{\mu}}_{g})^{2}$ (30)
$=:{\bm{\sigma}}_{g}^{2}/{\bm{\sigma}}^{2}$ (31)
Plugging this into Equation 25 gives:
$\text{GBN}_{\Delta}({\bm{X}})=\frac{{\bm{X}}-({\bm{\mu}}_{g}-{\bm{\mu}})}{{\bm{\sigma}}_{g}/{\bm{\sigma}}}$ (32)
Reducing the Train-Test Discrepancy
To mitigate the increased train-test discrepancy with smaller batch sizes in GBN, we propose to re-sample ghost batches for the calculation of normalization statistics. In Section 3.2 we described how we can replace Equation 21 with Equation 9, repeated below:
$\widetilde{\bm{X}}=\frac{\hat{\bm{X}}-\hat{{\mathbf{m}}}}{\hat{{\mathbf{s}}}}$ (33)
Where $\hat{{\mathbf{m}}}$ and $\hat{{\mathbf{s}}}$ are analogous to $\hat{{\bm{\mu}}}$ and $\hat{{\bm{\sigma}}}$, except they are computed over a random subset of $\hat{\bm{X}}$ for each element in the batch. By doing this, each sample only contributes weakly to its own normalization statistics in expectation, similar to batch normalization over the full accelerator batch instead of ghost batch normalization over a smaller ghost batch. We can extend this trick to $\text{GBN}_{\Delta}({\bm{X}})$ from Equation 32, replacing ${\bm{\mu}}_{g}$ and ${\bm{\sigma}}_{g}$ with $\mathbf{m}$} and $\mathbf{s}$}, computed like $\hat{{\mathbf{m}}}$ and $\hat{{\mathbf{s}}}$ except on the original inputs $\bm{X}$} instead of $\hat{{\bm{X}}}$. This gives the expression for GNI in Equation 10:
$\text{GNI}({\bm{X}})=\frac{\bm{X}-({\mathbf{m}}-\bm{\mu})}{{\mathbf{s}}/{\bm{\sigma}}}$ (34)
Appendix D Distribution of ghost noise
This part presents the calculation details of the noise distributions derived in Section 3.3. For $\hat{\mu}_{gc}$ we have:
$\operatorname{\operatorname{\mathbb{E}}}[\hat{\mu}_{gc}]=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nci}\right]=\frac{1}{NI}\sum_{n}\sum_{i}\operatorname{\operatorname{\mathbb{E}}}[x_{nci}]=0$ (35)
$\mathrm{Var}[\hat{\mu}_{gc}]$ $=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nci}\right]^{2}$ (36)
$=\operatorname{\operatorname{\mathbb{E}}}\left[\frac{1}{NI}\sum_{n}\sum_{i}x_{nci}-\mu_{nc}+\mu_{nc}\right]^{2}$ (37)
$=\frac{1}{N^{2}I^{2}}\sum_{n}\sum_{i}\sigma_{Ic}^{2}+\frac{1}{N^{2}I^{2}}NI^{2}\sigma_{Bc}^{2}$ (38)
$=\frac{1}{NI}\sigma_{Ic}^{2}+\frac{1}{N}\sigma_{Bc}^{2}$ (39)
For $\hat{\sigma}_{gc}^{2}$ we have:
$\hat{\sigma}_{gc}^{2}$ $=\frac{1}{NI}\sum_{n}\sum_{i}(x_{nci}-\mu_{nc}+\mu_{nc})^{2}$ (40)
$=\frac{1}{NI}\sum_{n}\sum_{i}\big((x_{nci}-\mu_{nc})^{2}+(\mu_{nc}-\mu)^{2}+2(x_{nci}-\mu_{nc})(\mu_{nc}-\mu)\big)$ (41)
@@ Appendix D Distribution of ghost noise
$\overset{\overline{x}_{nc}\approx\operatorname{\operatorname{\mathbb{E}}}x_{nc}}{\approx}\frac{1}{NI}\sum_{n}\sum_{i}\sigma_{Ic}^{2}\left(\frac{x_{nci}-\mu_{nc}}{\sigma_{Ic}}\right)^{2}+\frac{1}{N}\sigma_{Bc}^{2}\sum_{n}\left(\frac{\mu_{nc}-\mu}{\sigma_{Bc}}\right)^{2}+\frac{1}{N}(\mu_{nc}-\mu_{nc})(\mu_{nc}-\mu)$ (42)
$=\frac{1}{NI}\sum_{n}\sum_{i}\sigma_{Ic}^{2}\left(\frac{x_{nci}-\mu_{nc}}{\sigma_{Ic}}\right)^{2}+\frac{1}{N}\sigma_{Bc}^{2}\sum_{n}\left(\frac{\mu_{nc}-\mu}{\sigma_{Bc}}\right)^{2}$ (43)
Appendix E Additional Experimental Results
(a) Scale noise - epoch 1
(b) Shift noise - epoch 1
(c) Scale noise - epoch 100
(d) Shift noise - epoch 100
Figure 6: Measured scale and shift noise distributions from GNI versus our calculated analytical distribution in FNN
Distribution of Ghost Noise in FNNs
The ghost noise distribution for convolutional networks depends on the inter-sample and intra-sample variances which make it hard to model. Here we study the simpler 1d case, which does not have this effect and the ghost noise distribution primarily depends on the ghost batch size. We use a two-layer Fully connected neural network to classify MNIST (Deng 2012), with 512 and 300 neurons in each layer. The inputs are flattened image tensors. When the ghost batch size is 32, we plot out the distribution of measured scale and shift noise, with kernel density estimation to smooth the curve. We further plot out our calculated analytical distribution curve according to Section 3.3, i.e. analytical shift noise follows a normal distribution with mean 0 and std $\sqrt{1/32}$, and analytical scale noise distribution follows a Chi-square distribution with degree of freedom 32 and scale $1/32$. We show that our analytical solution is able to match the ghost noise distribution in 1d case.
Comparison to other Noise Injection Methods
In Section 5, we discussed two other noise injection methods. NoMorelization (Liu et al. 2022) is composed of two trainable scalars and a zero-centered Gaussian noise injector that is originally meant to replace the normalization layers (for use in noise-free settings). The explicit Gaussian Noise Injections proposed by Camuto et al. (2020) are also additive Gaussian Noise applied after every layer except the last one. We note that these methods are similar to AGNI-$\mu$, but the noise is sampled elementwise instead of channelwise. In Figure 7 we show the effect of applying i.i.d. Elementwise Additive Gaussian Noise (EAGN) of a fixed variance $\sigma$ (hyperparameter) in every normalization layer in the CIFAR-100 ResNet-18 setup from Figure 3. It is not fully clear how best to apply this noise so we consider two placements, either before or after the learned affine transformation. The results are shown in the figure, EAGN can provide a small accuracy boost but is unable to match GNI (which achieves over 78% accuracy here).
Figure 7: Extension of Figure 3 - Validation accuracy for EAGN
Applicability to MLPs
We trained an MLP (3 hidden layers of width 1024, each with batch norm + ReLU) on CIFAR-100 with a batch size of 1024, keeping the rest of the training setup the same as in our ResNet-18 experiments. After tuning the learning rate, the results are shown in Table 4. This suggests that GNI can also be beneficial in MLPs.
Baseline GNI-16 GNI-32